OpenStudy (anonymous):

For a metal? :) Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1 (1, 5) (-5, -1) (5, 1) (-5, 1)

OpenStudy (anonymous):

I think its 1,5 but im not sure

OpenStudy (akashdeepdeb):

Not quite. Which MeTal btw ? :') In the second equation! x = 1+4y [Now SUBSTITUTE this x value in the in the first equation] Like this! [ 2(1+4y) - 9y = 1 ] ---> Solve now! :D Understood? :)

OpenStudy (anonymous):

O.o i'm confused :P

OpenStudy (akashdeepdeb):

We need to SUBSTITUTE OVER HERE RIGHT? So we find any variables' value in terms of the other variable like we found x = 1+4y [In terms of y] So we SUBSTITUTE this value in the fiirst equation! 2(1+4y) - 9y = 1 2+8y-9y = 1 y=1 [Now SUBSTITUTE this value in any equation to FIND X] Let's put it in the second equation x = 1+4y = 1+4(1) = 5 So solution (x,y) = (5,1) Understood? :)

OpenStudy (anonymous):

So i was right though 5,1 is the answer?

OpenStudy (akashdeepdeb):

You said 1,5 buddy! Good job though! :) There is a goldmine of difference between (5,1) and (1,5) btw! :)

OpenStudy (anonymous):

Oh my mistake :P LOL Thanks sooo much!

OpenStudy (anonymous):

would you like another metal? I have another question